OpenStudy (anonymous):

Which point lies in the solution set for the following system of inequalities? y < 2x + 4 y < -2x + 2 (1, 0) (-5, -2) (0, -3) (-1, 5)

OpenStudy (anonymous):

Basically all you have to do is pluck each set of coords. in and see if the inequality stays true. Saying this, the 3rd answer is the only one that works. -3 < 2(0)+2= -3< 0+2= -3< 2 which is correct. Now we use the same coords. for the 2nd equation- -3< -2(0)+2 = -3< 0+2= -3 < 2

OpenStudy (anonymous):

that was right :)

OpenStudy (anonymous):

Lol yeah I know :) All you go to do with those is take the set of points which is (x,y) and put them in the inequality as x & y.

OpenStudy (anonymous):

If the set of points match both inequalities, then it's correct.

OpenStudy (anonymous):

Any other problem you need help with?

OpenStudy (anonymous):

yes lol one minute

OpenStudy (anonymous):

k.